# reference.bib

### v1

@@ V Conclusion
In this paper, we proposed EvRWKV, a novel and effective framework for low-light image enhancement that leverages continuous cross-modal interaction between event streams and low-light images through dual-domain processing. Our approach uniquely integrates the RWKV architecture with the EISFE module, enabling adaptive and robust fusion of asynchronous event data and standard images. The Cross-RWKV module effectively captures fine-grained temporal and spatial dependencies, while the EISFE module jointly performs noise suppression and structural restoration via frequency-domain filtering and spatial deformable convolutions. Extensive experiments on multiple challenging real-world low-light datasets demonstrate that EvRWKV significantly outperforms existing state-of-the-art methods in both quantitative metrics and visual quality, effectively suppressing noise, enhancing contrast, and restoring fine details even under severe degradation such as motion blur and extreme darkness. Our dual-domain fusion and adaptive gating mechanisms successfully balance complementary information from event and image modalities, resulting in superior enhancement robustness and generalization. In future work, we plan to explore further improvements in cross-modal fusion strategies and extend the application of the EvRWKV framework to other challenging imaging tasks such as real-time video enhancement and dynamic scene understanding in extreme lighting conditions.
[…]

### v2

@@ V Conclusion
In this paper, we proposed EvRWKV, a novel and effective framework for LLIE that leverages continuous cross-modal interaction between event streams and low-light images through dual-domain processing. Our approach uniquely integrates the RWKV architecture with the EISFE module, enabling adaptive and robust fusion of asynchronous event data and standard images. The Cross-RWKV module effectively captures fine-grained temporal and spatial dependencies, while the EISFE module jointly performs noise suppression and structural restoration via frequency-domain filtering and spatial deformable convolutions. Extensive experiments on multiple challenging real-world low-light datasets demonstrate that EvRWKV significantly outperforms existing state-of-the-art methods in both quantitative metrics and visual quality, effectively suppressing noise, enhancing contrast, and restoring fine details even under severe degradation such as motion blur and extreme darkness. Our dual-domain fusion and adaptive gating mechanisms successfully balance complementary information from event and image modalities, resulting in superior enhancement robustness and generalization. In future work, we plan to explore further improvements in cross-modal fusion strategies and extend the application of the EvRWKV framework to other challenging imaging tasks such as real-time video enhancement and dynamic scene understanding in extreme lighting conditions.